OpenStudy (anonymous):

1/2x+1/4x=16 A.16 B.21.3 C.32 D.64 WILL GIVE MEDAL FOR CORRECT ANSWER!!!!

OpenStudy (anonymous):

(1/2)x + (1/4)x = 16 2x + x = 64 3x = 64 x = 64/3 or 21 1/3

OpenStudy (anonymous):

Your answer is B, 21.3